OpenStudy (kidthatbro8):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax and x represents the original price of the item. : C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol.

OpenStudy (yanasidlinskiy):

What have you gotten so far?

OpenStudy (yanasidlinskiy):

What does X=?

OpenStudy (kidthatbro8):

nothing, yet. i don't really understand it. do i subtract 0.06 first ?

OpenStudy (yanasidlinskiy):

No. You do NOT subtract:) We are going to be multiplying since it asks how much:) What does X=? Do they not give it to you?

OpenStudy (kidthatbro8):

no, they don't, so i really have no idea.

OpenStudy (kidthatbro8):

oh, i think they say it was $100 up there.

OpenStudy (yanasidlinskiy):

Lol, so X=100:) Lets do 1st step:) C = 0.765(100) + 0.06(0.765*100) Keep on doing it plugging everything in. Once you finish let me know:)

OpenStudy (kidthatbro8):

c = 76.5 + 0.06(76.5) ?

OpenStudy (yanasidlinskiy):

Yep! There you go!!:) Sorry, I had to go to the RS, lol!!

OpenStudy (kidthatbro8):

okay c: do you multiply 0.06(76.5) now ?

OpenStudy (yanasidlinskiy):

Actually you add. Do you know what numbers? Looks like you did something wrong on your side. I have different numbers

OpenStudy (kidthatbro8):

i guess not. i'm really confused about this ? what did you do ?

OpenStudy (yanasidlinskiy):

Oh wait!! Ok. I see what you did. It's actually on my side!! I though asked something different. Yes. You'd multiply and THEN you add!! Sorry about the confusion!!

OpenStudy (kidthatbro8):

oh, okay ! so it would be c = 76.5 + 4.59 ?

OpenStudy (yanasidlinskiy):

Yes! Correct! You're getting it!!! Yay!!!

OpenStudy (kidthatbro8):

c = 81.09 ??? :v he still paid allot for the coat tho

OpenStudy (yanasidlinskiy):

Hahha!!! Yes!! That'd be your answer!!! I agree!! There are coats though that are a lot more expensive. I think it depends off the material thoughxD Sorry for replying late.